OpenStudy (anonymous):

Find the sum of 2√6 and √24 . i dont understand how to solve this , please help

OpenStudy (lgbasallote):

\[2\sqrt 6 + \sqrt{24} \implies 2 \sqrt 6 + \sqrt{4 \times 6} \implies 2\sqrt 6 + 2\sqrt 6 \] can you solve that now?

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

yes i understand now , appreciate it

OpenStudy (lgbasallote):

\[\huge \color{purple}{\textbf{<tips hat>}}\]